# Perturbative soft graviton theorems in de Sitter spacetime

### v2

@@ 6.2 Subleading and sub-subleading corrections
In this section, we determine the subleading and sub-subleading corrections to the flat space soft graviton theorem and comment on its universality. We use the asymmetric form of the scalar field propagator (3.3) for the ease of the calculations. Let us substitute the soft graviton mode (35), scalar field mode (10), and the scalar field propagator (3.3) into eq. (54) and separate the $\mathcal{O}(l^{-2})$ terms as follows:
$\Gamma_{n+1}^{(2)}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}2\kappa\int d^{4}zd^{4}z^{\prime}\int\dfrac{d^{4}p}{(2\pi)^{4}}\bigg{(}\dfrac{4m^{4}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{4}}+\dfrac{2m^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{3}}+\dfrac{2(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{2}}\\ -\dfrac{p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta}}{m^{2}(p^{2}+m^{2})}+\dfrac{9i(p_{i}\cdot z^{\prime})(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{8m^{2}(p^{2}+m^{2})}+\dfrac{15(p_{i}\cdot z^{\prime})^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{8m^{2}(p^{2}+m^{2})}-\dfrac{13i(p_{i}\cdot z^{\prime})^{3}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{12m^{2}(p^{2}+m^{2})}\\ -\dfrac{(p_{i}\cdot z^{\prime})^{4}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{8m^{2}(p^{2}+m^{2})}-\dfrac{2im^{2}(p\cdot z)(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{3}}-\dfrac{i(p\cdot z)(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{2}}+\dfrac{m^{2}z^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})^{2}}+\dfrac{z^{\prime 2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{4(p^{2}+m^{2})}\\ -\dfrac{iz^{\prime 2}(p_{i}\cdot z^{\prime})(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{p^{2}+m^{2}}-\dfrac{z^{\prime 2}(p_{i}\cdot z^{\prime})^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{4(p^{2}+m^{2})}-\dfrac{m^{2}z^{\prime 4}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{8(p^{2}+m^{2})}-\dfrac{z^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}+\dfrac{i(p^{\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{p^{2}+m^{2}}\bigg{)}\\ \cdot e^{i(p-p_{i}-k)\cdot z^{\prime}}e^{-ip\cdot z}\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}.$ (68)
The terms inside the parentheses in the above expression can be simplified in a similar manner as we did in the previous section. First, we replace $z^{\prime\mu}$ by $-ie^{-ip\cdot z^{\prime}}\frac{\partial}{\partial p_{\mu}}e^{ip\cdot z^{\prime}}$, and then do integration by parts in $p$ to move exponential factor $e^{ip\cdot z^{\prime}}$ outside the derivative operator $\frac{\partial}{\partial p_{\mu}}$. We then perform the $z^{\prime}$-integral on the exponential factor $e^{i(p-p_{i}-k)\cdot z^{\prime}}$ to generate the delta function $\delta^{(4)}(p-p_{i}-k)$. Finally, we perform $p$-integral on the delta function and take the soft limit.
We evaluated each term inside the parentheses in (68) in a similar manner as described above and obtained the following expression [^5] We must note that from eq. (55) the structure of $\Gamma_{n+1}^{(2)}$ is as follows: $\Gamma_{n+1}^{(2)}=A^{(2)}\Gamma_{n}^{(0)}+A^{(1)}\Gamma_{n}^{(1)}+A^{(0)}\Gamma_{n}^{(2)}$. After evaluating and simplifying eq. (68), one needs to identify and separate the above terms to reach at eq. (69).:
$\Gamma_{n+1}^{(2)}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\kappa\int d^{4}z\bigg{(}-\dfrac{m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{3}}+\dfrac{im^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{2(p_{i}\cdot k)^{3}}+\dfrac{m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)^{2}}{(p_{i}\cdot k)^{3}}\\ -\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p_{i}\cdot k)^{2}}-\dfrac{im^{2}(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{(p_{i}\cdot k)^{2}}-\dfrac{2m^{2}(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{(p_{i}\cdot k)^{2}}+\dfrac{m^{2}(z^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{p_{i}\cdot k}\bigg{)}e^{-ip_{i}\cdot z}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}e^{-ip_{j}\cdot z_{j}}(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}+A^{(1)}\Gamma_{n}^{(1)}(\{p_{i}\})+A^{(0)}\Gamma_{n}^{(2)}(\{p_{i}\}),$ (69)
where $\mathcal{O}(l^{-2})$ corrections to the flat space soft factor are
$A^{(2)}=-m^{2}\sum_{i=1}^{n}\kappa\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot k)^{3}}+\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{m^{2}(p_{i}\cdot k)^{2}}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot k)^{3}}k\cdot\partial_{p_{i}}-\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{(p_{i}\cdot k)^{2}}\partial_{p_{i}}^{\beta}\bigg{)}+\mathcal{O}(k^{-1}).$ (70)
We note that $A^{(2)}$ includes two types of terms: $1/k^{3}$ and $1/k^{2}$, which obviously do not contribute at the same order. Thus, as in (56), we expand the soft factor in $\delta$ and $\omega$ to separate the subleading and sub-subleading order corrections to the flat space soft factor in the above expression, which are as follows:
$S^{(2,-1)}=-m^{2}\sum_{i=1}^{n}\kappa\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot\hat{k})^{3}},$ (71)
@@
$S^{(2,0)}=-m^{2}\sum_{i=1}^{n}\kappa\bigg{(}\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{m^{2}(p_{i}\cdot\hat{k})^{2}}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot\hat{k})^{3}}\hat{k}\cdot\partial_{p_{i}}-\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{(p_{i}\cdot\hat{k})^{2}}\partial_{p_{i}}^{\beta}\bigg{)}.$ (72)
Rest of the coefficients $S^{(r,s)}$ in eq. (56) vanish for the tree level scattering processes.
Consistency checks:

### v3

@@ 6.2 Subleading and sub-subleading corrections
In this section, we determine the subleading and sub-subleading corrections to the flat space soft graviton theorem. We use the asymmetric form of the scalar field propagator (3.3) for the ease of the calculations. Let us substitute the soft graviton modes (39), scalar field modes (10), and the scalar field propagator (3.3) into (61) and separate the $\mathcal{O}(l^{-2})$ terms as follows:
@@
$\Gamma_{n+1}^{(2)}(\{p_{1},...,p_{n}\},k)=-i\sum_{i=1}^{n}\kappa\int d^{4}zd^{4}z^{\prime}\int\dfrac{d^{4}p}{(2\pi)^{4}}\bigg(\dfrac{4m^{4}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{4}}+\dfrac{2m^{2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{3}}$
$+\dfrac{2(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{2}}-\dfrac{17p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{8m^{2}(p^{2}+m^{2})}+\dfrac{39i(p_{i}\cdot z^{\prime})(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{8m^{2}(p^{2}+m^{2})}+\dfrac{41(p_{i}\cdot z^{\prime})^{2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{8m^{2}(p^{2}+m^{2})}$
$-\dfrac{19i(p_{i}\cdot z^{\prime})^{3}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{12m^{2}(p^{2}+m^{2})}-\dfrac{(p_{i}\cdot z^{\prime})^{4}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{8m^{2}(p^{2}+m^{2})}+\dfrac{3z^{\prime 2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}-\dfrac{3iz^{\prime 2}(p_{i}\cdot z^{\prime})(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}$
$-\dfrac{z^{\prime 2}(p_{i}\cdot z^{\prime})^{2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{4(p^{2}+m^{2})}-\dfrac{m^{2}z^{\prime 4}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{8(p^{2}+m^{2})}-\dfrac{2im^{2}(p\cdot z)(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{3}}-\dfrac{i(p\cdot z)(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{2}}$
$+\dfrac{m^{2}z^{2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})^{2}}-\dfrac{z^{2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}+\dfrac{3i(p^{\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}+\dfrac{5(p_{i}\cdot z^{\prime})(p^{\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}-\dfrac{i(p_{i}\cdot z^{\prime})^{2}(p^{\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}$
$-\dfrac{im^{2}z^{\prime 2}(p^{\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}+\dfrac{(p\cdot z^{\prime})(p_{i}^{\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}-\dfrac{(p\cdot p_{i})(z^{\prime\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{4(p^{2}+m^{2})}\bigg)e^{i(p-p_{i}-k)\cdot z^{\prime}}$
$\cdot e^{-ip\cdot z}\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle.$ (75)
The terms inside the parentheses in the above expression can be simplified in a similar manner as we did for $\mathcal{O}(l^{-1})$ terms in (64) in the previous section. In brief, we first replace all $z^{\prime}$ by $i\partial_{p}$, and then do integration by parts in $p$ to move exponential factor $e^{ip\cdot z^{\prime}}$ outside the derivative operator $\partial_{p}$. We then combine all the exponential factors to $e^{i(p-p_{i}-k)\cdot z^{\prime}}$ and perform the $z^{\prime}$-integral to generate the delta function $\delta^{(4)}(p-p_{i}-k)$. Finally, we perform $p$-integral on the delta function and take the soft limit.
We evaluated each term inside the parentheses in (6.2) in a similar manner as described above and obtained the following expression:
$\Gamma_{n+1}^{(2)}(\{p_{1},...,p_{n}\},k)=-i\sum_{i=1}^{n}\frac{\kappa}{2}\int d^{4}z\bigg(\dfrac{3m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{3}}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot k)^{2}}-\dfrac{17(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{8m^{2}(p_{i}\cdot k)}$
$+\dfrac{im^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{2(p_{i}\cdot k)^{3}}-\dfrac{2i(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{(p_{i}\cdot k)^{2}}+\dfrac{m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)^{2}}{(p_{i}\cdot k)^{3}}-\dfrac{3i(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)}{2(p_{i}\cdot k)^{2}}$
$+\dfrac{39i(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)}{8m^{2}(p_{i}\cdot k)}-\dfrac{2(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)(k\cdot z)}{(p_{i}\cdot k)^{2}}+\dfrac{41(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)^{2}}{8m^{2}(p_{i}\cdot k)}+\dfrac{i(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)^{2}(k\cdot z)}{2(p_{i}\cdot k)^{2}}$
$-\dfrac{19i(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)^{3}}{12m^{2}(p_{i}\cdot k)}-\dfrac{(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)^{4}}{8m^{2}(p_{i}\cdot k)}+\dfrac{z^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{p_{i}\cdot k}+\dfrac{im^{2}z^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{2(p_{i}\cdot k)^{2}}$
$-\dfrac{3iz^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)}{2(p_{i}\cdot k)}-\dfrac{z^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)^{2}}{4(p_{i}\cdot k)}-\dfrac{m^{2}z^{4}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{8(p_{i}\cdot k)}-\dfrac{im^{2}(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{2}}+\dfrac{4i(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{p_{i}\cdot k}$
$-\dfrac{2m^{2}(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{(p_{i}\cdot k)^{2}}+\dfrac{(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{2(p_{i}\cdot k)}+\dfrac{5(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)}{p_{i}\cdot k}-\dfrac{i(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)^{2}}{p_{i}\cdot k}$
$-\dfrac{im^{2}z^{2}(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{p_{i}\cdot k}+\dfrac{5m^{2}(z^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{4(p_{i}\cdot k)}-\dfrac{z^{\alpha}z^{\beta}\varepsilon_{\alpha\beta}}{4}\bigg)\bigg(1-ik\cdot z-\dfrac{1}{2}(k\cdot z)^{2}+...\bigg)e^{-ip_{i}\cdot z}$
$\cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle,$ (76)
which can be rewritten in the desired form as [^7] We must note that from (62), the structure of $\Gamma_{n+1}^{(2)}$ is as follows: $\Gamma_{n+1}^{(2)}=A^{(2)}\Gamma_{n}^{(0)}+A^{(1)}\Gamma_{n}^{(1)}+A^{(0)}\Gamma_{n}^{(2)}$. To put Eq. (6.2) in the above desired form, one needs to identify the terms corresponding to $A^{(0)}$ and $A^{(1)}$ in (6.2) and separate them from the remaining terms, which correspond to $A^{(2)}$, to reach at the final expression (6.2).
$\Gamma_{n+1}^{(2)}(\{p_{1},...,p_{n}\},k)=-i\sum_{i=1}^{n}\frac{\kappa}{2}\int d^{4}z\bigg(\dfrac{3m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{3}}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot k)^{2}}-\dfrac{im^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{(p_{i}\cdot k)^{3}}$
$-\dfrac{3i(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)}{2(p_{i}\cdot k)^{2}}-\dfrac{im^{2}(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{2}}+\mathcal{O}(k^{-1})\bigg)e^{-ip_{i}\cdot z}$
$\cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}e^{-ip_{j}\cdot z_{j}}(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle+A^{(1)}\Gamma_{n}^{(1)}+A^{(0)}\Gamma_{n}^{(2)},$ (77)
where $\mathcal{O}(k^{-1})$ terms contribute beyond the sub-subleading order and are suppressed. To transform the soft factor in the momentum space, we replace all $z$ by $i\partial_{p_{i}}$ in the above expression. Thus, the $\mathcal{O}(l^{-2})$ correction to the flat space soft factor takes the following form:
$A^{(2)}=\sum_{i=1}^{n}\frac{\kappa}{2}\bigg($ $\dfrac{3m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{3}}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot k)^{2}}+\dfrac{m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p_{i}\cdot k)^{3}}k\cdot\partial_{p_{i}}+\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{2}}p_{i}\cdot\partial_{p_{i}}$
$+\dfrac{m^{2}(p_{i}^{\alpha}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{2}}\partial_{p_{i}}^{\beta}\bigg)+\mathcal{O}(k^{-1}).$ (78)
We note that $A^{(2)}$ includes two types of terms: $1/k^{3}$ and $1/k^{2}$, which obviously do not contribute at the same order. Thus, as in (63), we expand the soft factor in $\delta$ to separate the subleading and sub-subleading order corrections to the flat space soft factor in the above expression. The subleading and sub-subleading corrections take the following forms, respectively.
$S^{(2,-1)}=\sum_{i=1}^{n}\frac{\kappa}{2}\dfrac{3m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2\omega(p_{i}\cdot\hat{k})^{3}},$ (79)
$S^{(2,0)}=\sum_{i=1}^{n}\frac{\kappa}{2}\bigg(\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot\hat{k})^{2}}+\dfrac{m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p_{i}\cdot\hat{k})^{3}}\hat{k}\cdot\partial_{p_{i}}+\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot\hat{k})^{2}}p_{i}\cdot\partial_{p_{i}}+\dfrac{m^{2}(p_{i}^{\alpha}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot\hat{k})^{2}}\partial_{p_{i}}^{\beta}\bigg).$ (80)
In the following section, we evaluate the Feynman diagram where a soft graviton is attached to an internal line.
6.3 Graviton attached to an internal line
Figure 3: Soft graviton attached to an internal line.
Let us now consider the Feynman diagram in which the graviton is attached to an internal line (see Fig. (3)), which produces the subleading soft graviton theorem in the flat space. We need to check whether such a diagram will introduce additional perturbative corrections to the flat space soft factor in de Sitter background. If the $\mathcal{S}$-matrix for this scattering process is defined by $N_{n+1}(\{p_{1},...,p_{n}\},k)$, then it can be decomposed in the soft expansion as follows:
[…]
Since the soft graviton is not attached to an external line, this diagram does not produce $\frac{1}{\omega}$-pole terms. Thus, the above expression can be expanded at $\omega\to 0$ as follows:
$N_{n+1}(\{p_{1},...,p_{n}\},k)=N^{(0)}_{n+1}(\{p_{1},...,p_{n}\},0)+\delta N^{(1)}_{n+1}(\{p_{1},...,p_{n}\},0)\\ +\delta^{2}N^{(2)}_{n+1}(\{p_{1},...,p_{n}\},0)+$..., (82)
where $N^{(0)}_{n+1}$ produce the flat space subleading soft factor, $N^{(1)}_{n+1}$ and $N^{(2)}_{n+1}$ are correction terms. Thus, the total $\mathcal{S}$-matrix, after including these additional terms to Eq. (61), is
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\frac{\kappa_{i}}{2}\int d^{4}zd^{4}z^{\prime}\left(1-\dfrac{z^{2}}{l^{2}}-\dfrac{z^{\prime 2}}{l^{2}}\right)f^{*\rho\sigma}(z^{\prime},k)\bigg\{2\partial^{\prime}_{\rho}g^{*}_{p_{i}}(z^{\prime})\partial^{\prime}_{\sigma}D(z^{\prime},z)\bigg\}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle+N_{n+1}(\{p_{1},...,p_{n}\},k).$ (83)
The $\mathcal{S}$-matrix for the scattering process depicted in Fig. (2) has already been computed in Section (6.1). Thus, we will not re-do these calculations, and simply add the additional correction terms to Eq. (72). The resultant expression is
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\frac{\kappa}{2}\bigg(\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}k\cdot\partial_{p_{i}}-\dfrac{im}{l}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot k)^{2}}k\cdot\partial_{p_{i}}+\dfrac{2mi}{l}\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\partial_{p_{i}}^{\beta}\bigg)\Gamma^{(0)}_{n}\\ +\dfrac{1}{l}A^{(0)}\Gamma^{(1)}_{n}+\varepsilon_{\alpha\beta}N^{(0)\alpha\beta}_{n+1}(\{p_{1},...,p_{n}\},0)+\delta\varepsilon_{\alpha\beta}N^{(1)\alpha\beta}_{n+1}(\{p_{1},...,p_{n}\},0),$ (84)
where the second term arises from the soft expansion of the exponential factor $e^{-ik\cdot z^{\prime}}$, and $N_{n+1}$ is decomposed as $\varepsilon_{\alpha\beta}N_{n+1}^{\alpha\beta}$. The gauge invariance allows us to evaluate additional correction terms $N^{(0)\alpha\beta}_{n+1}$ and $N^{(1)\alpha\beta}_{n+1}$. The gauge transformation does not receive perturbative corrections at this order, i.e., $\varepsilon_{\alpha\beta}\to\varepsilon_{\alpha\beta}-ik_{\alpha}\xi_{\beta}-ik_{\beta}\xi_{\alpha}$. Thus, under this gauge transformation, the $\mathcal{S}$-matrix should remain invariant, i.e.,
$\sum_{i=1}^{n}\frac{\kappa}{2}\bigg(2(p_{i}\cdot\xi)+2(p_{i}\cdot\xi)k\cdot\partial_{p_{i}}+\dfrac{2mi}{l}\xi_{\beta}\partial^{\beta}_{p_{i}}\bigg)\Gamma^{(0)}_{n}+2k_{\alpha}\xi_{\beta}N^{(0)\alpha\beta}_{n+1}+\dfrac{2}{\omega l}k_{\alpha}\xi_{\beta}N^{(1)\alpha\beta}_{n+1}=0,$ (85)
where the first term vanishes due to the momentum conservation. Equating the remaining terms order by order, we obtain [^8] After deducing $N_{n+1}^{(0)\alpha\beta}$ and $N_{n+1}^{(1)\alpha\beta}$ from Eq. (85), we have put them in the symmetrized form in Eqs. (86) and (87).
$N_{n+1}^{(0)\alpha\beta}=-\dfrac{1}{2}\sum_{i=1}^{n}\dfrac{\kappa}{2}\big(p_{i}^{\alpha}\partial^{\beta}_{p_{i}}+p_{i}^{\beta}\partial^{\alpha}_{p_{i}}\big)\Gamma_{n}^{(0)},$ (86)
$N_{n+1}^{(1)\alpha\beta}=-\dfrac{im}{2}\sum_{i=1}^{n}\dfrac{\kappa}{2}\bigg(\dfrac{p_{i}^{\alpha}}{p_{i}\cdot\hat{k}}\partial^{\beta}_{p_{i}}+\dfrac{p_{i}^{\beta}}{p_{i}\cdot\hat{k}}\partial^{\alpha}_{p_{i}}\bigg)\Gamma_{n}^{(0)}.$ (87)
These are the additional contributions to the flat space soft graviton theorem from the diagram where soft graviton is attached to an internal line. After plugging these corrections, Eqs. (86) and (87), into Eq. (84), we obtained the final expression:
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\frac{\kappa}{2}\bigg(\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}-i\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}k_{\gamma}J^{\beta\gamma}}{p_{i}\cdot k}-\dfrac{m}{l}\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}k_{\gamma}J^{\beta\gamma}}{(p_{i}\cdot k)^{2}}\bigg)\Gamma^{(0)}_{n}+\dfrac{1}{l}A^{(0)}\Gamma^{(1)}_{n}$, (88)
where $J^{\beta\gamma}=i\big(p_{i}^{\beta}\partial^{\gamma}_{p_{i}}-p_{i}^{\gamma}\partial^{\beta}_{p_{i}}\big)$ is the angular momentum operator in the momentum space. The second term in the above expression is the flat space subleading soft graviton theorem, and the third term is the leading correction to the flat space soft factor.
$S^{(1,0)}=-m\sum_{i=1}^{n}\frac{\kappa}{2}\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}\hat{k}_{\gamma}J^{\beta\gamma}}{(p_{i}\cdot\hat{k})^{2}}$ (89)
We must note that the subleading corrections to the flat space soft factor do not receive additional contributions from the diagram where soft graviton is attached to an internal line. It is because the $N_{n+1}^{(2)}$ correction term arises at order $\mathcal{O}(\delta^{2})$ in the soft expansion. Thus, the subleading corrections (79) remain the same even after including the contributions from the diagram in Fig. (3). To check whether the sub-subleading corrections (80) do not receive contributions from $N_{n+1}^{(2)}$, we need to analyze the gauge invariance of them explicitly in the momentum space, which is extremely complicated and we will not attempt to do it in the present work.
6.4 Consistency checks